OpenStudy (anonymous):

2x + y = 7 5x + y = 9 Solve the system of equations.

OpenStudy (anonymous):

(-2/3, 25/3) (2/3, 17/3) (2/3, 25/3)

OpenStudy (anonymous):

multiply the first one by \(-1\) all the way across to get \[-2x-y=-7\\ 5x+y=9\] then add in a column and the y terms will go bye bye

rishavraj (rishavraj):

http://www.purplemath.com/modules/systlin5.htm

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you got this or no?

OpenStudy (anonymous):

No..lol

rishavraj (rishavraj):

u need to subtract both equations..... u will get -3x = -2 x = 2/3

OpenStudy (anonymous):

Okay... What about y?

rishavraj (rishavraj):

plug x = 2/3 in ant equation and get "y"... hey u should visit tht link

OpenStudy (anonymous):

I'm not clicking on any link that some random person posts... It's probably some type of virus or scam

rishavraj (rishavraj):

lol...its not tht....:) bt whatever ....

OpenStudy (anonymous):

How am I supposed to know it's not?

rishavraj (rishavraj):

hey forget it....don't visit......

OpenStudy (anonymous):

Lol I definitely won't.

rishavraj (rishavraj):

lol as u wish....

OpenStudy (anonymous):

Mmhmm.